# MEMoE: Enhancing Model Editing with Mixture of Experts Adaptors

### v1

@@ Abstract
Model editing aims to efficiently alter the behavior of Large Language Models (LLMs) within a desired scope, while ensuring no adverse impact on other inputs. Recent years have witnessed various model editing methods been proposed. However, these methods either exhibit poor overall performance or struggle to strike a balance between generalization and locality. We propose MOMoE, a model editing adapter utilizing a Mixture of Experts (MoE) architecture with a knowledge anchor routing strategy. MOMoE updates knowledge using a bypass MoE structure, keeping the original parameters unchanged to preserve the general ability of LLMs. And, the knowledge anchor routing ensures that inputs requiring similar knowledge are routed to the same expert, thereby enhancing the generalization of the updated knowledge. Experimental results show the superiority of our approach over both batch editing and sequential batch editing tasks, exhibiting exceptional overall performance alongside outstanding balance between generalization and locality. Our code will be available.
@@ 1 Introduction
In light of these, we propose MOMoE, a novel framework leveraging the MoE architecture alongside knowledge anchor routing to enhance the overall performance of model editing. Considering the intrinsic sparsity of knowledge information and the advantage of MoE in handling sparse features [3, 45], MOMoE extends upon a parallel MoE structure to enhance the accuracy of knowledge learning. This MoE-style adapter is confined to only one layer of the model, preserving all original parameters, thereby enhancing the locality of model editing and further reducing the impact on model’s general ability. On the other hand, prior research has highlighted the considerable benefits stemming from the specialize feature of MoE’s experts in multi-task learning, and pointed out that appropriate routing strategy can lead to improved generalization performance [57, 50]. We introduce the knowledge anchor routing strategy, enabling routers to selectively focus on specific knowledge aspects of the inputs, ensuring that queries requiring similar knowledge are routed to the same experts. Through this “professional people do professional things” approach, the generalization performance of MOMoE has been improved.
@@
  • We present MOMoE, a method utilizing a MoE architecture with a knowledge anchor routing strategy for model editing.
  • Experiments show that our proposed method achieves state-of-the-art editing performance. MOMoE achieves high accuracy while effectively balancing generality and locality.
@@
Figure 1: Left: Apart from our MOMoE, no method achieves both high accuracy and high balance. Right: Significant room for improvement in the overall performance of current methods.
@@ 2 Preliminaries of Model Editing
Figure 2: The architecture of MOMoE, compared with conventional MoE. Same color denote inputs requiring same knowledge. Pentagrams symbolize knowledge anchors within the input sentences, while squares and triangles represent ordinary input tokens during editing process and generality evaluation respectively. The distribution of tokens within the FFN illustrates that knowledge anchor consolidate inputs requiring same knowledge to the same experts.
@@ 3 Methodology
In this section, we provide a detailed introduction to MOMoE, a model editing adapter based on MoE structure and knowledge anchor routing strategy, as shown in Figure 2. This method achieves a balance between generality and locality, while enabling highly precise model editing.
3.1 MOMoE Architecture
One of the core ideas of MOMoE is to introduce several MOE-style experts via bypasses to facilitate knowledge updates and learning, while freezing all the original parameters of LLM to maintain its general ability to the greatest extent. The right of Figure 2 illustrates the forward process of MOMoE, sharing both similarities and distinctions with traditional MoE depicted on the left.
Similar to traditional MoE, MOMoE employs a structure that integrates multiple parallel experts within the transformer feed-forward network (FFN). The choice to use the FFN module is not only due to its traditional role in MoE [20] but also aligns with recent experimental findings of knowledge probing technologies that the MLP layers within FFN store knowledge [7, 33, 34].
Differently, MOMoE incorporate the additional parallel experts through a bypass mechanism, thereby preserving the original parameters of the model to enhance the locality of model editing. This bypass structure also provides potential to further enhance the generalization performance of model editing (more details in §3.2). Moreover, this adaptation is applied to only one layer of the model. The selective addition of adapters in one layer is based on two considerations: first, previous model editing techniques have demonstrated that modifying parameters in a single layer can effectively achieve knowledge updates [33, 34]; second, this strategy further maintains the structure and parameters of the original model, ensuring its general ability is preserved to the greatest extent.
Specifically, given token $x_{i}$ in the input sequence $X=\{x_{i}\}_{i=1}^{L}$, MOMoE with $E$ experts first introduces a gate decision vector $\mathcal{G}\in\mathbb{R}^{E}$ that dispatches different input tokens to different experts, which is calculated as:
@@ 3.1 MOMoE Architecture
Overall, the forward process of the MOMoE layer, combined with the frozen original parameters $\mathbf{W}_{0}$, can be expressed as:
@@ 3.2 Knowledge Anchor Routing
Another core idea of MOMoE is the routing strategy based on the knowledge anchors. Inspired by the specialize nature of experts in MoE architecture [53], we aim to route inputs requiring similar knowledge to the same expert during both training and testing phases, thereby enhancing the model’s generalization performance when dealing with new knowledge.
In MOMoE, we define the named entities in input sentences as “knowledge anchors”. For example, in the input “Who is the president of the United States?” the entities “president” and “United States” serve as knowledge anchors. The routing strategy allocate tokens to the appropriate experts based on these anchors, ensuring that inputs requiring similar knowledge are routed to the same expert. The effect is demonstrated in Figure 2, showing the token distribution within the FFN. This approach better captures and retains the semantic associations of knowledge in input data. Consequently, it enhances the model’s generalization performance when handling knowledge and also optimizes the efficiency of expert utilization to a certain extent (as validated in §5.2).
@@ 4 Experiments
In this section, we first describe our experimental setup. Then, we show the remarkable performance of MOMoE on two challenging model editing tasks: batch editing and sequential batch editing.
@@ 4.1 Experimental Setups
We select GPT2-XL and LLaMA2-7B as the base models. We opted for the more challenging model editing tasks: batch editing and sequential batch editing, to evaluate the performance of MOMoE. For batch editing, following [29], the batch size is set to 30 and the model is rolled back to the initial state after each batch editing. For sequential batch editing, the batch size is 10 for a total of 1000 edits, without rollback. Further details of the baselines and the implementation are provided in the Appendix D.
@@ 4.2 Batch Editing
We first evaluate the effectiveness of MOMoE under batch editing settings. The evaluation results are presented in Table 1. For all models and all metrics, our method consistently achieves the best scores. MOMoE’s reliability scores are all above 90, generalization scores are all above 85, and locality scores are perfect at 100. The improvements across various metrics are significant. Compared to GPT2-XL, our method demonstrates even more remarkable improvements on LLaMA2-7B, with a maximum improvement of up to 17.55 points in accuracy and 56.21 in generality. In the §5.2, we conduct further experimental analysis to explore the reasons behind the substantial enhancement in generalization. Considering some current researches concern that model editing methods may significantly affect a model’s general ability [14, 15, 40], we perform a more detailed general ability evaluation using a broader task datasets in §5.1.
@@ 4.3 Sequential Batch Editing
We evaluate MOMoE on 1,000 samples from both datasets for sequential batch editing. The evaluation is conducted after the entire editing process is completed. The results, shown in Table 2, indicate that MOMoE achieves the best scores in most cases. It only ranks second because GRACE [16] achieves perfect scores of 100 in some matrices. Similar to the batch editing results, MOMoE performs better on LLaMA2-7B, demonstrating a significant advantage in generality while maintaining accuracy and locality much close to 100. Regarding GRACE, it excels in reliability and locality but performs poorly in generality due to its use of a codebook to memorize encountered editing instances [16, 12]. However, its poor performance in generality suggests a problem with regurgitation. Overall, the reliability and generality consistently lag behind in comparison to batch editing, indicating there is still room for improvement in this field.
@@ 5 Detailed Analysis and Discussion
In this section, we conduct a further evaluation and analysis of the performance of MOMoE. Firstly, we assess the impact of MOMoE on the model’s general ability using a broader range of task datasets. Secondly, we show potential sources of MOMoE’s generalization advantage through an analysis of expert specialize phenomena. Finally, we present an extensive serious of ablation study to evaluate the efficacy of various model configurations, including the number of experts, the target layer and the routing strategies.
@@ 5.1 General Ability Test
Figure 3: Performance on general tasks of edited models using MOMoE, MEMIT and MEND, with different batch sizes for edits.
@@
We conduct evaluations on LLaMA2-7B based on batch editing settings, progressively increasing the batch size to show the impact of more edited samples (the model is rolled back to the initial state after each batch editing). The results are shown in the Figure 3. It is important to note that any increase or decrease in performance metrics implies an impact on model’s general ability. Compared to the MEMIT and MEND, the MOMoE yields consistently stable model performance under various batch editing conditions. With the increase in batch size and edited samples, both MEMIT and MEND significantly diminish the model’s general ability, while the influence of MOMoE fluctuates within a smaller range. This further corroborates MOMoE’s advantage in locality score in §4.2.
@@ 5.2 Experts Specialize
In order to fully analyze the behavior of experts routing when dealing with different knowledge inputs, we first conduct a static analysis based the trained MOMoE module in 4.2. That is to say, at this stage, the editing data is fixed; we analyze the behavior of experts solely through the model inference process. Further, we dynamically select the knowledge to be edited as input to observe the behavior of knowledge anchor routing. The results are shown in Table 3. As the concentration of the input knowledge categories increases, the consistency of the experts also improves, and the generality simultaneously rise. Additionally, the consistency of the same experts handling the same knowledge is higher and closely aligned with the generalization scores. Given the near-perfect accuracy score, we speculate that errors in generalization assessments may be due to incorrect routing of inputs to the wrong expert. Analysis of the bad cases in the generalization test supports this hypothesis. By adhering to the principle that “professional people do professional things”, the strategy of routing inputs requiring similar or same knowledge to the same expert proves effective in improving knowledge generalization.
@@ 5.3 Ablation Study
How does the number of experts impact the performance of MOMoE? The left plot in Figure 4 illustrates the performance of MOMoE with different numbers of experts. Due to computational resource limitations, we could only add up to 6 additional experts. Aligning with the primary experimental results in §4, we set the $top_{k}$ value to 1. We find that the reliability and locality of model editing do not change with the number of experts; there is neither a decrease nor an improvement in performance. However, the generalization of knowledge fluctuated with the number of experts, peaking when the number of experts is 4. Inspired by [53], we hypothesize that this optimal experts number is related to editing batch size. When the number of editing samples is not large, more experts may introduce interference, thereby reducing the generalization performance.
@@
What is the optimal layer for applying MOMoE? As shown in the middle of Figure 4, similar to the results of experts number experiment, the reliability and locality score remain unaffected, with only the generality score exhibiting fluctuations and peaking at the 16th layer. The best editing layer identified from these validation experiments aligns with the results obtained using knowledge probes technology in [33]. Combining the findings from above experiments, we can infer that the accuracy and locality performance are inherently guaranteed by the characteristics of the bypass MoE structure, which is consistent with the design principles of MOMoE discussed in §3.1.
@@
What is the best routing strategy in MOMoE? In Figure 4, the rightmost plot illustrates the performance when using various routing strategies for MOMoE. Specifically, we compare the soft merging [53] of experts with discrete top-1, top-2, and top-3 routing strategy. The top-1 routing setting yields the best performance. Further, based on the experiment introduced in §5.2, we examined the consistency score of routing under different values of $k$. We observed that as k increases, more experts participate in the computation, but the consistency score decreases. We believe that this inconsistency in expert utilization leads to the decrease in generalization performance. Further experimental results, detailed in the Appendix E, demonstrate that regardless of the expert number and the target layer, the top-1 performance remains the best. In addition, the discrete top-1 routing has an advantage in computational efficiency by requiring only one experts to be activated during inference.
@@
From Table 1 and 2, it is evident that MOMoE demonstrates a significant performance advantage in batch editing tasks over sequential editing. To further evaluate MOMoE’s ability in batch editing, we progressively increased the batch size. Experimental results are shown in Table 4. When the batch size for batch editing reached 1000, equivalent to the total number of sequential batch editing, MOMoE exhibited significant performance advantages. Reliability and locality are scarcely affected by the increase in batch size, maintaining close to 100. Meanwhile, the generality score surpassed by 33.07 points, highlighting MOMoE’s performance advantage in batch editing. As for the decline in sequential editing, our analysis of bad cases indicates that can be attributed to catastrophic forgetting: edits complete earlier are more prone to errors.
@@ 6 Conclusion
In this paper, we present MOMoE, a model editing adapter utilizing MoE architecture with knowledge anchor routing strategy. MOMoE updates knowledge using a bypass MoE structure, keeping the original parameters unchanged to preserve the model’s general ability. And, the knowledge anchor ensures that questions requiring similar knowledge are handled by the same expert, thereby enhancing the generalization of the updated knowledge. Experiment results demonstrate that our method significantly surpasses all compared baselines. MOMoE shows near-perfect accuracy and locality scores close to 100, along with a generalization score exceeding 90, indicating exciting promise for practical applications of model editing technology.

### v2

@@ Abstract
Model editing aims to efficiently alter the behavior of Large Language Models (LLMs) within a desired scope, while ensuring no adverse impact on other inputs. Recent years have witnessed various model editing methods been proposed. However, these methods either exhibit poor overall performance or struggle to strike a balance between generalization and locality. We propose MEMoE, a model editing adapter utilizing a Mixture of Experts (MoE) architecture with a knowledge anchor routing strategy. MEMoE updates knowledge using a bypass MoE structure, keeping the original parameters unchanged to preserve the general ability of LLMs. And, the knowledge anchor routing ensures that inputs requiring similar knowledge are routed to the same expert, thereby enhancing the generalization of the updated knowledge. Experimental results show the superiority of our approach over both batch editing and sequential batch editing tasks, exhibiting exceptional overall performance alongside outstanding balance between generalization and locality. Our code will be available.
@@ 1 Introduction
In light of these, we propose MEMoE, a novel framework leveraging the MoE architecture alongside knowledge anchor routing to enhance the overall performance of model editing. Considering the intrinsic sparsity of knowledge information and the advantage of MoE in handling sparse features [3, 45], MEMoE extends upon a parallel MoE structure to enhance the accuracy of knowledge learning. This MoE-style adapter is confined to only one layer of the model, preserving all original parameters, thereby enhancing the locality of model editing and further reducing the impact on model’s general ability. On the other hand, prior research has highlighted the considerable benefits stemming from the specialize feature of MoE’s experts in multi-task learning, and pointed out that appropriate routing strategy can lead to improved generalization performance [57, 50]. We introduce the knowledge anchor routing strategy, enabling routers to selectively focus on specific knowledge aspects of the inputs, ensuring that queries requiring similar knowledge are routed to the same experts. Through this “professional people do professional things” approach, the generalization performance of MEMoE has been improved.
@@
  • We present MEMoE, a method utilizing a MoE architecture with a knowledge anchor routing strategy for model editing.
  • Experiments show that our proposed method achieves state-of-the-art editing performance. MEMoE achieves high accuracy while effectively balancing generality and locality.
@@
Figure 1: Left: Apart from our MEMoE, no method achieves both high accuracy and high balance. Right: Significant room for improvement in the overall performance of current methods.
@@ 2 Preliminaries of Model Editing
Figure 2: The architecture of MEMoE, compared with conventional MoE. Same color denote inputs requiring same knowledge. Pentagrams symbolize knowledge anchors within the input sentences, while squares and triangles represent ordinary input tokens during editing process and generality evaluation respectively. The distribution of tokens within the FFN illustrates that knowledge anchor consolidate inputs requiring same knowledge to the same experts.
@@ 3 Methodology
In this section, we provide a detailed introduction to MEMoE, a model editing adapter based on MoE structure and knowledge anchor routing strategy, as shown in Figure 2. This method achieves a balance between generality and locality, while enabling highly precise model editing.
3.1 MEMoE Architecture
One of the core ideas of MEMoE is to introduce several MOE-style experts via bypasses to facilitate knowledge updates and learning, while freezing all the original parameters of LLM to maintain its general ability to the greatest extent. The right of Figure 2 illustrates the forward process of MEMoE, sharing both similarities and distinctions with traditional MoE depicted on the left.
Similar to traditional MoE, MEMoE employs a structure that integrates multiple parallel experts within the transformer feed-forward network (FFN). The choice to use the FFN module is not only due to its traditional role in MoE [20] but also aligns with recent experimental findings of knowledge probing technologies that the MLP layers within FFN store knowledge [7, 33, 34].
Differently, MEMoE incorporate the additional parallel experts through a bypass mechanism, thereby preserving the original parameters of the model to enhance the locality of model editing. This bypass structure also provides potential to further enhance the generalization performance of model editing (more details in §3.2). Moreover, this adaptation is applied to only one layer of the model. The selective addition of adapters in one layer is based on two considerations: first, previous model editing techniques have demonstrated that modifying parameters in a single layer can effectively achieve knowledge updates [33, 34]; second, this strategy further maintains the structure and parameters of the original model, ensuring its general ability is preserved to the greatest extent.
Specifically, given token $x_{i}$ in the input sequence $X=\{x_{i}\}_{i=1}^{L}$, MEMoE with $E$ experts first introduces a gate decision vector $\mathcal{G}\in\mathbb{R}^{E}$ that dispatches different input tokens to different experts, which is calculated as:
@@ 3.1 MEMoE Architecture
Overall, the forward process of the MEMoE layer, combined with the frozen original parameters $\mathbf{W}_{0}$, can be expressed as:
@@ 3.2 Knowledge Anchor Routing
Another core idea of MEMoE is the routing strategy based on the knowledge anchors. Inspired by the specialize nature of experts in MoE architecture [53], we aim to route inputs requiring similar knowledge to the same expert during both training and testing phases, thereby enhancing the model’s generalization performance when dealing with new knowledge.
In MEMoE, we define the named entities in input sentences as “knowledge anchors”. For example, in the input “Who is the president of the United States?” the entities “president” and “United States” serve as knowledge anchors. The routing strategy allocate tokens to the appropriate experts based on these anchors, ensuring that inputs requiring similar knowledge are routed to the same expert. The effect is demonstrated in Figure 2, showing the token distribution within the FFN. This approach better captures and retains the semantic associations of knowledge in input data. Consequently, it enhances the model’s generalization performance when handling knowledge and also optimizes the efficiency of expert utilization to a certain extent (as validated in §5.2).
@@ 4 Experiments
In this section, we first describe our experimental setup. Then, we show the remarkable performance of MEMoE on two challenging model editing tasks: batch editing and sequential batch editing.
@@ 4.1 Experimental Setups
We select GPT2-XL and LLaMA2-7B as the base models. We opted for the more challenging model editing tasks: batch editing and sequential batch editing, to evaluate the performance of MEMoE. For batch editing, following [29], the batch size is set to 30 and the model is rolled back to the initial state after each batch editing. For sequential batch editing, the batch size is 10 for a total of 1000 edits, without rollback. Further details of the baselines and the implementation are provided in the Appendix D.
@@ 4.2 Batch Editing
We first evaluate the effectiveness of MEMoE under batch editing settings. The evaluation results are presented in Table 1. For all models and all metrics, our method consistently achieves the best scores. MEMoE’s reliability scores are all above 90, generalization scores are all above 85, and locality scores are perfect at 100. The improvements across various metrics are significant. Compared to GPT2-XL, our method demonstrates even more remarkable improvements on LLaMA2-7B, with a maximum improvement of up to 17.55 points in accuracy and 56.21 in generality. In the §5.2, we conduct further experimental analysis to explore the reasons behind the substantial enhancement in generalization. Considering some current researches concern that model editing methods may significantly affect a model’s general ability [14, 15, 40], we perform a more detailed general ability evaluation using a broader task datasets in §5.1.
@@ 4.3 Sequential Batch Editing
We evaluate MEMoE on 1,000 samples from both datasets for sequential batch editing. The evaluation is conducted after the entire editing process is completed. The results, shown in Table 2, indicate that MEMoE achieves the best scores in most cases. It only ranks second because GRACE [16] achieves perfect scores of 100 in some matrices. Similar to the batch editing results, MEMoE performs better on LLaMA2-7B, demonstrating a significant advantage in generality while maintaining accuracy and locality much close to 100. Regarding GRACE, it excels in reliability and locality but performs poorly in generality due to its use of a codebook to memorize encountered editing instances [16, 12]. However, its poor performance in generality suggests a problem with regurgitation. Overall, the reliability and generality consistently lag behind in comparison to batch editing, indicating there is still room for improvement in this field.
@@ 5 Detailed Analysis and Discussion
In this section, we conduct a further evaluation and analysis of the performance of MEMoE. Firstly, we assess the impact of MEMoE on the model’s general ability using a broader range of task datasets. Secondly, we show potential sources of MEMoE’s generalization advantage through an analysis of expert specialize phenomena. Finally, we present an extensive serious of ablation study to evaluate the efficacy of various model configurations, including the number of experts, the target layer and the routing strategies.
@@ 5.1 General Ability Test
Figure 3: Performance on general tasks of edited models using MEMoE, MEMIT and MEND, with different batch sizes for edits.
@@
We conduct evaluations on LLaMA2-7B based on batch editing settings, progressively increasing the batch size to show the impact of more edited samples (the model is rolled back to the initial state after each batch editing). The results are shown in the Figure 3. It is important to note that any increase or decrease in performance metrics implies an impact on model’s general ability. Compared to the MEMIT and MEND, the MEMoE yields consistently stable model performance under various batch editing conditions. With the increase in batch size and edited samples, both MEMIT and MEND significantly diminish the model’s general ability, while the influence of MEMoE fluctuates within a smaller range. This further corroborates MEMoE’s advantage in locality score in §4.2.
@@ 5.2 Experts Specialize
In order to fully analyze the behavior of experts routing when dealing with different knowledge inputs, we first conduct a static analysis based the trained MEMoE module in 4.2. That is to say, at this stage, the editing data is fixed; we analyze the behavior of experts solely through the model inference process. Further, we dynamically select the knowledge to be edited as input to observe the behavior of knowledge anchor routing. The results are shown in Table 3. As the concentration of the input knowledge categories increases, the consistency of the experts also improves, and the generality simultaneously rise. Additionally, the consistency of the same experts handling the same knowledge is higher and closely aligned with the generalization scores. Given the near-perfect accuracy score, we speculate that errors in generalization assessments may be due to incorrect routing of inputs to the wrong expert. Analysis of the bad cases in the generalization test supports this hypothesis. By adhering to the principle that “professional people do professional things”, the strategy of routing inputs requiring similar or same knowledge to the same expert proves effective in improving knowledge generalization.
@@ 5.3 Ablation Study
How does the number of experts impact the performance of MEMoE? The left plot in Figure 4 illustrates the performance of MEMoE with different numbers of experts. Due to computational resource limitations, we could only add up to 6 additional experts. Aligning with the primary experimental results in §4, we set the $top_{k}$ value to 1. We find that the reliability and locality of model editing do not change with the number of experts; there is neither a decrease nor an improvement in performance. However, the generalization of knowledge fluctuated with the number of experts, peaking when the number of experts is 4. Inspired by [53], we hypothesize that this optimal experts number is related to editing batch size. When the number of editing samples is not large, more experts may introduce interference, thereby reducing the generalization performance.
@@
What is the optimal layer for applying MEMoE? As shown in the middle of Figure 4, similar to the results of experts number experiment, the reliability and locality score remain unaffected, with only the generality score exhibiting fluctuations and peaking at the 16th layer. The best editing layer identified from these validation experiments aligns with the results obtained using knowledge probes technology in [33]. Combining the findings from above experiments, we can infer that the accuracy and locality performance are inherently guaranteed by the characteristics of the bypass MoE structure, which is consistent with the design principles of MEMoE discussed in §3.1.
@@
What is the best routing strategy in MEMoE? In Figure 4, the rightmost plot illustrates the performance when using various routing strategies for MEMoE. Specifically, we compare the soft merging [53] of experts with discrete top-1, top-2, and top-3 routing strategy. The top-1 routing setting yields the best performance. Further, based on the experiment introduced in §5.2, we examined the consistency score of routing under different values of $k$. We observed that as k increases, more experts participate in the computation, but the consistency score decreases. We believe that this inconsistency in expert utilization leads to the decrease in generalization performance. Further experimental results, detailed in the Appendix E, demonstrate that regardless of the expert number and the target layer, the top-1 performance remains the best. In addition, the discrete top-1 routing has an advantage in computational efficiency by requiring only one experts to be activated during inference.
@@
From Table 1 and 2, it is evident that MEMoE demonstrates a significant performance advantage in batch editing tasks over sequential editing. To further evaluate MEMoE’s ability in batch editing, we progressively increased the batch size. Experimental results are shown in Table 4. When the batch size for batch editing reached 1000, equivalent to the total number of sequential batch editing, MEMoE exhibited significant performance advantages. Reliability and locality are scarcely affected by the increase in batch size, maintaining close to 100. Meanwhile, the generality score surpassed by 33.07 points, highlighting MEMoE’s performance advantage in batch editing. As for the decline in sequential editing, our analysis of bad cases indicates that can be attributed to catastrophic forgetting: edits complete earlier are more prone to errors.
@@ 6 Conclusion
In this paper, we present MEMoE, a model editing adapter utilizing MoE architecture with knowledge anchor routing strategy. MEMoE updates knowledge using a bypass MoE structure, keeping the original parameters unchanged to preserve the model’s general ability. And, the knowledge anchor ensures that questions requiring similar knowledge are handled by the same expert, thereby enhancing the generalization of the updated knowledge. Experiment results demonstrate that our method significantly surpasses all compared baselines. MEMoE shows near-perfect accuracy and locality scores close to 100, along with a generalization score exceeding 90, indicating exciting promise for practical applications of model editing technology.